Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR3 in R61 @ 14:00 pm on January 10th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • SS Magnet status
  • Hydrogen status
  • Issues arising from previous running
  • Shutdown work in Jan and Feb
  • Operations :
    • Shifts open from Feb 13 to March 6. Signup is slow. Please sign up. Run plans are in development.
    • Controls system check from cold to data-taking : Jan 23 - 24 (Ajit/Ed)
    • Expert-led run : Feb 10 - 14
  • AOB

Minutes

  • SS Magnet status (JBoehm)
    All 3 CC magnets are cold and full. All still in solenoid mode. Will put FC into flip and reverse SSU.
    The magnets have not run since they tripped last year. Should run and check for "damage" soon.
    Heater performance on FC is worse - pressure went up by 10 mbar - with the A/C down.
    1. Ramp FC alone to 165 A(?) in flip mode. FC group to define this threshold.
    2. Then test FC + SSU to confirm forces
    3. Extended test of full channel to wait until ops period when we have shift cover

    JBoehm has link box wiring documentation near-complete.
    Windows updates on QP system are now disabled, but need to remove one of the ones that sneaked itself on [PF].

  • Hydrogen status (VBayliss)
    Cooldown over Christmas: absorber reached 20K and cooler 12 K.
    12 K would freeze LH2; don't have heater control circuits in R9.
    Will do test with Neon in the next couple of weeks to confirm that a liquid would condense and drip into absorber pot, reducing the temperature difference between them.
  • Issues arising from previous running
    A complete list of the issues encountered during last data-taking period needs to be built up.
    Need to make sure that they will actually be dealt with (or at least ignored consciously).
    • e.g. What specific actions led to the PPS search drop?
    • e.g. Why did miceraid5 run out of disk space, when mitigation actions were decided in Feb. 2016? Have they actually been implemented yet?
    • Need to check eLog/paper log/MOM notebook and find what else is lurking

    Need the list for next week's meeting, as we are running out of time in which to make fixes and test
    PHodgson [as DC] will ask Melissa for copies of her lists.

  • Shutdown work in Jan and Feb
    The official version
    The Beamstop and Potterax door have both been serviced. HN thinks that, with both chillers active, it should be possible to run conventional magnets at full power.
    After some discussion, the present plan is:
    • Wed. 11th Jan. Test FC in flip mode, and conventional magnet heat load test.
    • Thu. 12th Jan. FC + SSU test, including diffuser check.
    • Fri. 13th Jan. SSD M2 50A test.
    • 16th-20th John Govans plumbing
    • 23 & 24th John Govans plumbing and Controls system check (Ajit/Ed)
    • Wed. 25th? PPS functional test - HALL CLOSED

    The ODH system should be serviced now if possible, rather than waiting until Easter.
    Also need target BPS test and MICEnet lockout #1399.

  • Operations :
    • Shifts open from Feb 13 to March 6. Signup is slow. Please sign up. Run plans are in development.
    • Controls system check from cold to data-taking : Jan 23 - 24 (Ajit/Ed) - do NOT power cycle DL Hardware IOCs
    • Expert-led run : Feb 10 - 14 (Fri - Mon!)
  • AOB
  • What is status of neutron monitor débâcle?
    Ours is still with ISIS Health Physics - we should get an update at Monday's MICE-ISIS Safety Mtg.
    The temporary monitor is still running OK but we should get the calibration checked.

Diffuser_field_test.txt Magnifier - Diffuser field test? (1.15 KB) Cobb, John, 10 January 2017 11:49